what is the quadratic equation for the given solutions 7i and -7i
Answer by stanbon(75887)

what is the quadratic equation for the given solutions 7i and -7i
----
Ans: (x-7i)(x--7i) = x^2+7ix-7ix+49 = x^2+49
--------------------------------
Cheers,
Stan H.